### v1

@@ 2.4 Uniform $(n;k)$ model
Throughout this paper, we use examples of specific multi-district school choice problems to demonstrate particular phenomena, some of which stand in contrast to the propositions in Pathak and Sönmez (2008) that hold for the single-district setting. To analyze how frequently such phenomena occur, we consider large random two-district school choice problems inspired by the uniform models of Babaioff et al. (2019). In the uniform $(n;k)$ model, there are $2$ districts labelled $L$ and $R$ [^6] Later, we denote schools in $L$ as $\ell_{1},\ell_{2},...$ and schools in $R$ as $r_{1},r_{2},...$ for ease of reading. We use $L$ and $R$ rather than numbering the districts $1$ and $2$ for this reason. Collectively, $L$ and $R$ contain $n$ students $I=\{i_{1},...,i_{n}\}$ and $n$ schools $S=\{s_{1},s_{2},...,s_{n}\}$, each with unit capacity (i.e., $q_{s}=1$ for all $s\in S$).
Each student is either sincere or sophisticated; is either constrained or unconstrained; and resides in either district $L$ or district $R$. There are thus eight categories of students: One for each possible sophistication type - constraint type - district of residence combination. A student’s category is drawn independently of all other students’ categories, and there is a positive probability of a student’s category being any of the eight possibilities. As such, there exists some $p>0$ such that for each category, the probability of an arbitrary student being in this category is at least $p$. Each student’s preference ordering over schools (which may include schools in any district) is drawn uniformly at random from among all (strict) possible preference orderings of length $k$ [^7] This is a special case of the procedure used to draw preference lists in Immorlica and Mahdian (2005). Each student’s preference ordering is independent of all other students’ preference orderings, and of all students’ categories.
Each school independently has probability $\nicefrac{{1}}{{2}}$ of being located in district $L$ and $\nicefrac{{1}}{{2}}$ of being located in district $R$ [^8] These probabilities need only be constant and nonzero values that sum to $1$ for our results to hold, but we set them equal to avoid clutter. Finally, each school has a complete (strict) priority ordering over all students, drawn uniformly at random from the set of all such possible orderings, and independently of everything else. Thus, for any school and any two students $i_{a}$ and $i_{b}$, the probability that $i_{a}$ has priority over $i_{b}$ at that school is $\nicefrac{{1}}{{2}}$
@@ 3 Sophistication Types
In this section, we show that it is possible for a sophisticated student to prefer that (i.e., strictly benefit if) some sincere student becomes sophisticated. In fact, we prove that such students are abundant in large random markets. This result stands in contrast to Pathak and Sönmez (2008), which shows that in the single-district setting, all sophisticated students weakly suffer if any sincere student becomes sophisticated.
@@ 3.1 Example: A sophisticated student may prefer for a sincere student to become sophisticated
We first provide an illustrative example. Suppose that there are two districts with schools $\ell_{1},\ell_{2}\in L$ and school $r_{1},r_{2},r_{3},r_{4}\in R$, where each school has unit capacity. District $L$ uses DA, while district $R$ uses BM. Further suppose that there are five students $i_{1}$, $i_{2}$, $i_{3}$, $i_{4}$, and $i_{5}$.
The students’ preference orderings, districts of residence, constraint types, and sophistication types are as follows. [^9] The notation $s_{a}\succ s_{b}$ indicates a preference for school $s_{a}$ over school $s_{b}$. Students whose preference orderings contain schools in only one district reside in that district and have arbitrary constraint types. Other students (those whose preference orderings contains schools in both districts) are unconstrained and reside in an arbitrary district.
[…]
The schools’ priority orderings include the following. [^10] The notation $i_{a}-i_{b}$ indicates priority for student $i_{a}$ over student $i_{b}$. Technically, a school’s priority ordering must include all students. Here, for each school, we list only the priority ordering over students who find the school acceptable, as no other student would ever apply to the school. Other students could be placed anywhere in each school’s priority ordering without affecting our results.
@@
$\ell_{2}:$ $i_{2}-i_{3}$
$r_{1}:$ $i_{2}-i_{5}$
$r_{2}:$ $i_{4}-i_{2}$
$r_{3}:$ $i_{3}$
$r_{4}:$ $i_{5}$
We will show that $i_{3}$ prefers for $i_{2}$ to become sophisticated, and that it is irrelevant which definition of sincere-unconstrained is used.
First, consider the original setting where $i_{2}$ is sincere. Observe that regardless of which definition of sincere-unconstrained is used, $i_{2}$ will enroll in district $L$. If $i_{2}$ does not strategize at all, then $i_{2}$ will enroll in district $L$ because her first choice school is in district $L$. If $i_{2}$ strategizes over districts but always reports a truthful ROL, then $i_{2}$ will see that enrolling in district $R$ results in her being unassigned. This is because $i_{2}$ would not be matched in the first round of BM, during which both $r_{1}$ and $r_{2}$ would be filled. Because $L$ uses DA, on the other hand, $i_{2}$ would not be matched in the first round but would still get matched with a school in her preference ordering. Therefore, $i_{2}$ enrolls in district $L$.
Since $i_{2}$ enrolls in district $L$, sophisticated-unconstrained student $i_{3}$ enrolls in district $R$ (and ranks only $r_{3}$) to avoid being unmatched. The matching process results in $i_{1}$ assigned to $\ell_{1}$, $i_{2}$ assigned to $\ell_{2}$, $i_{3}$ assigned to $r_{3}$, $i_{4}$ assigned to $r_{2}$, and $i_{5}$ assigned to $r_{1}$. This is the unique Nash equilibrium outcome. Note that in this outcome, $i_{2}$ is assigned to $i_{2}$’s fourth-choice school and $i_{3}$ is assigned to $i_{3}$’s second-choice school.
Suppose instead that $i_{2}$ becomes sophisticated. Student $i_{2}$ has no chance of being assigned to her first- or second-choice schools, and is guaranteed admittance to $r_{1}$ if she enrolls in district $R$ and ranks $r_{1}$ first, so she does so. Note that if $i_{2}$ does not rank $r_{1}$ first, then $i_{2}$ would not be admitted to $r_{1}$. Student $i_{3}$ therefore chooses to enroll in district $L$ (and ranks only $l_{2}$), as this guarantees her admittance at $\ell_{2}$ (and she would not be admitted there otherwise). The matching process results in $i_{1}$ assigned to $\ell_{1}$, $i_{2}$ assigned to $r_{1}$, $i_{3}$ assigned to $\ell_{2}$, $i_{4}$ assigned to $r_{2}$, and $i_{5}$ is assigned to $r_{4}$. This is the unique Nash equilibrium outcome. In this outcome, $i_{2}$ is assigned to $i_{2}$’s third-choice school and $i_{3}$ is assigned to $i_{3}$’s first-choice school. Therefore, $i_{3}$ prefers for $i_{2}$ to become sophisticated.
@@ 3.2 Large-Market Analysis
We generalize the example from Section 3.1 to the uniform $(n;4)$ model. (The same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 4$.) We say that a sophisticated student $i_{a}$ prefers for a sincere student $i_{b}$ to become sophisticated if $i_{a}$ strictly prefers her match in every Nash equilibrium of the multi-district choice problem when $i_{b}$ is sophisticated to her match in every Nash equilibrium of the multi-district choice problem when $i_{b}$ is sincere. We show that there can be many sophisticated students who prefer for distinct sincere students to become sophisticated.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;4)$ model with one district using DA and the other using BM, there exists a set of sophisticated students of expected size at least $\tau n$ where each sophisticated student strictly prefers for a distinct sincere student to become sophisticated, and weakly prefers for all other sincere students to become sophisticated.
@@
WLOG, let $L$ be the district that uses DA and let $R$ be the one that uses BM. We start by lower bounding the expected number of ordered quintuples of students $(i_{1},i_{2},i_{3},i_{4},i_{5})$ that satisfy the following conditions (as in the example from Section 3.1):
@@
    Conditions on sophistication types, constraint types, and districts of residence:
@@
      $i_{3}$ is sophisticated and unconstrained.
    4. (d)
      Both $i_{4}$ and $i_{5}$ are sincere and reside in $R$.
@@
      $i_{1}$ most prefers a school $\ell_{1}\in L$.
@@
      $i_{5}$ most prefers a school $r_{1}\in R$ and second most prefers a school $r_{4}\in R$.
@@
      $i_{4}$ most prefers a school $r_{2}\in R$.
@@
      $i_{3}$ most prefers a school $\ell_{2}\in L$ and second most prefers a school $r_{3}\in R$.
@@
      $i_{2}$ has preference ordering $\ell_{1}\succ r_{2}\succ r_{1}\succ\ell_{2}$.
    6. (f)
      No other student finds $\ell_{1},\ell_{2},r_{1},r_{2},r_{3}$, or $r_{4}$ acceptable.
@@
      $i_{2}$ has priority over $i_{5}$ at $r_{1}$.
    4. (d)
      $i_{4}$ has priority over $i_{2}$ at $r_{2}$.
@@
As in the example in Section 3.1, in this situation, $i_{3}$ would strictly benefit from $i_{2}$ becoming sophisticated.
@@
We want to determine the probability that this set of conditions occurs for a specific $(i_{1},i_{2},i_{3},i_{4},i_{5})$. Each of the three sets of conditions is independent. The conditions on student sophistication types, constraint types, and districts of residence are satisfied with probability at least $p^{5}$. The conditions on student preferences and school locations are satisfied with probability
$\left(\frac{1}{2}\right)\cdot\left(\frac{n-1}{n}\cdot\frac{1}{2}\cdot\frac{n-2}{n-1}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-3}{n}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-4}{n}\cdot\frac{1}{2}\cdot\frac{n-5}{n-1}\cdot\frac{1}{2}\right)\cdot\\ \cdot\left(\frac{1}{n(n-1)(n-2)(n-3)}\right)\cdot\left(\frac{n-6}{n}\cdot\frac{n-7}{n-1}\cdot\frac{n-8}{n-2}\cdot\frac{n-9}{n-3}\right)^{n-5}.$
@@
In the expression above, each of the six parenthetical expressions corresponds to one of the six conditions on student preferences and school locations, and represents the condition’s probability, conditioned on the previous conditions. For sufficiently large $n$, [^12] Note that $\lim_{n\rightarrow\infty}(1-\frac{r}{n-a})^{(n-b)}$, where $a$, $b$, and $r$ are constants, is $e^{-r}$. this probability can be lower bounded by [^13] The final $\nicefrac{{1}}{{2}}$ is to accommodate for all the multiplicands of the form $\frac{n-a}{n-b}$ or $(1-\frac{r}{n-a})^{(n-b)}$, which have a nonzero limit, not fully converging for finite large $n$.
$\frac{1}{2^{6}}\cdot\frac{(n-4)!}{n!}\cdot\frac{1}{e^{24}}\cdot\frac{1}{2}.$
@@
Finally, the conditions on school priorities are satisfied with probability $\nicefrac{{1}}{{2^{4}}}$.
@@
The number of ordered quintuples of distinct students is $\tfrac{n!}{(n-5)!}$. List all such quintuples and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{I_{j}}$ be the indicator random variable that represents whether the students in $I_{j}$ satisfy all of the above conditions. Then the expected number of quintuples of students that satisfy all conditions is
$\mathbb{E}\left[\sum_{j=1}^{\tfrac{n!}{(n-5)!}}\mathbbm{1}_{I_{j}}\right]\geq\frac{n!}{(n-5)!}\cdot p^{5}\cdot\frac{1}{2^{4}}\cdot\frac{1}{2^{7}\cdot e^{24}}\cdot\frac{(n-4)!}{n!}=\frac{p^{5}(n-4)}{2^{11}\cdot e^{24}}.$
@@
To conclude the proof, we observe that by construction, no student can participate in two different quintuples that satisfy the above conditions. Therefore, all such quintuples of students involve distinct students, so there are at least an expected $\frac{p^{5}}{2^{11}\cdot e^{24}}\cdot(n-4)$ sophisticated students who each strictly prefer for a distinct sincere student to become sophisticated. Also by construction, each such sophisticated student is unaffected by sincere students outside of her quintuple, and the sophistication types of students $i_{1}$, $i_{4}$, and $i_{5}$ in the quintuple does not affect the dynamic between students $i_{3}$ and $i_{2}$. [^14] This is because $i_{1}$ and $i_{4}$ will always be assigned to their first choice schools, and $i_{5}$ will either be assigned to $r_{1}$ if $i_{2}$ is sincere or $r_{4}$ if $i_{2}$ is sophisticated. $r_{4}$ is part of this example to preclude $i_{5}$ from affecting the other students in the quintuple if she were to become sophisticated. Thus, these sophisticated students also weakly prefer for all other sincere students to become sophisticated.
We choose [^15] We made no attempt to optimize this value, as our goal was only to ascertain linearity in $n$ of the number of such students. $\tau=\frac{p^{5}}{2^{12}\cdot e^{24}}$ to satisfy the theorem statement. ∎

### v2

@@ 2.4 Uniform $(n;k)$ model
Throughout this paper, we use examples of specific multi-district school choice problems to demonstrate particular phenomena, some of which stand in contrast to the propositions in Pathak and Sönmez (2008) that hold for the single-district setting. To analyze how frequently such phenomena occur, we consider large random two-district school choice problems inspired by the uniform models of Babaioff et al. (2019). In the uniform $(n;k)$ model, there are $2$ districts labelled $L$ and RR. [^6] Later, we denote schools in $L$ as $\ell_{1},\ell_{2},$... and schools in $R$ as $r_{1},r_{2},$... for ease of reading. Collectively, $L$ and $R$ contain $n$ students $I=\{i_{1},...,i_{n}\}$ and $n$ schools $S=\{s_{1},s_{2},...,s_{n}\}$, each with unit capacity (i.e., $q_{s}=1$ for all $s\in S$).
Each student is either sincere or sophisticated; is either constrained or unconstrained; and resides in either district $L$ or district $R$. There are thus eight categories of students: one for each possible sophistication type – constraint type – district of residence combination. A student’s category is drawn independently of all other students’ categories, and there is a positive probability of a student’s category being any of the eight possibilities. As such, there exists some category-probability lower bound $p>0$, such that for each category, the probability of an arbitrary student being in this category is at least $p$. Each student’s preference ordering over schools (which may include schools in any district) is drawn uniformly at random from among all possible (strict) preference orderings of length kk. [^7] This is a special case of the procedure used to draw preference lists in Immorlica and Mahdian (2005). Each student’s preference ordering is independent of all other students’ preference orderings, and of all students’ categories.
Each school independently has probability $\nicefrac{{1}}{{2}}$ of being located in district $L$ and $\nicefrac{{1}}{{2}}$ of being located in district RR. [^8] These probabilities need only be constant and nonzero (and sum to $1$) for our results to hold, but we set them equal to avoid clutter. Finally, each school has a complete (strict) priority ordering over all students, drawn uniformly at random from the set of all such possible orderings, and independently of everything else. Thus, for any school and any two students $i_{a}$ and $i_{b}$, the probability that $i_{a}$ has priority over $i_{b}$ at that school is $\nicefrac{{1}}{{2}}$.
@@ 3 Sophistication Types
In this section, we show that it is possible for a sophisticated student to prefer that (i.e., strictly benefit if) some sincere student becomes sophisticated. In fact, we prove that such students are abundant in large random markets in which at least one district uses BM. This result stands in contrast to Proposition 3 of Pathak and Sönmez (2008), which is that in the single-district setting when BM is used, all sophisticated students weakly suffer if any sincere student becomes sophisticated. Finally, we prove that this phenomenon cannot occur when all districts use DA.
@@ 3.1 Example: A sophisticated student may prefer for a sincere student to become sophisticated
We first provide an illustrative example. Suppose that there are two districts, $L$ and $R$, with schools $\ell_{1}\in L$ and $r_{1},r_{2},r_{3},r_{4}\in R$ respectively, where each school has unit capacity. District $L$ uses an arbitrary mechanism, while district $R$ uses BM. Further suppose that there are four students, $i_{1}$, $i_{2}$, $i_{3}$, and $i_{4}$.
The students’ sophistication types and preference orderings, as well as the schools’ priority orderings, are shown in the table below. [^9] The notation $s_{a}\succ s_{b}$ indicates a student preference for school $s_{a}$ over school $s_{b}$. The notation $i_{a}-i_{b}$ indicates a school priority for student $i_{a}$ over student $i_{b}$. Technically, a school’s priority ordering must include all students. Here, for each school, we list only the priority ordering over students who find the school acceptable, as no other student would ever apply to the school. Other students could be placed anywhere in each school’s priority ordering without affecting our results. Students whose preference orderings contain schools in only one district reside in that district and have arbitrary constraint types. Students whose preference orderings contains schools in both districts are unconstrained and reside in an arbitrary district.
Student preferences
$\text{(sincere) }i_{1}:$ $r_{2}\succ r_{1}\succ\ell_{1}$
$\text{(sophisticated) }i_{2}:$ $\ell_{1}\succ r_{3}$
$\text{(sincere) }i_{3}:$ $r_{2}$
$\text{(sincere) }i_{4}:$ $r_{1}\succ r_{4}$
School priority orderings
@@
$r_{1}:$ $i_{1}-i_{4}$
$r_{2}:$ $i_{3}-i_{1}$
$r_{3}:$ $i_{2}$
$r_{4}:$ $i_{4}$
We show that $i_{2}$ strictly prefers for $i_{1}$ to become sophisticated. First, consider the original setting where $i_{1}$ is sincere-unconstrained, and recall that $i_{1}$ strategizes over districts but always reports a truthful ROL. Observe that $i_{1}$ enrolls in district $L$, as doing so matches her with school $\ell_{1}$ while enrolling in district $R$ results in her being unassigned. This is because if $i_{1}$ were to enroll in $R$, she would not be matched in the first round of BM, during which both $r_{1}$ and $r_{2}$ would be filled.
Since $i_{1}$ enrolls in district $L$, sophisticated-unconstrained student $i_{2}$ enrolls in district $R$ (and ranks only $r_{3}$) to avoid being unassigned. The matching process results in $i_{1}$ assigned to $\ell_{1}$, $i_{2}$ assigned to $r_{3}$, $i_{3}$ assigned to $r_{2}$, and $i_{4}$ assigned to $r_{1}$. This is the unique Nash equilibrium outcome. Note that in this outcome, $i_{2}$ is assigned to her second-choice school.
Suppose instead that $i_{1}$ becomes sophisticated. Student $i_{1}$ has no chance of being assigned to her first-choice school, and is guaranteed admittance to $r_{1}$ if and only if she enrolls in district $R$ and ranks $r_{1}$ first, so she does so. Student $i_{2}$ therefore chooses to enroll in district $L$ (and ranks only $\ell_{1}$), as this (and only this) guarantees her admittance at $\ell_{1}$. The matching process results in $i_{1}$ assigned to $r_{1}$, $i_{2}$ assigned to $\ell_{1}$, $i_{3}$ assigned to $r_{2}$, and $i_{4}$ is assigned to $r_{4}$. This is the unique Nash equilibrium outcome. In this outcome, $i_{2}$ is assigned to her first-choice school, which is a strict improvement for her compared to when $i_{1}$ is sincere.
@@ 3.2 Large-Market Analysis
We generalize the example from Section 3.1 to the uniform $(n;3)$ model. (The same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 3$.) We say that a sophisticated student $i_{a}$ strictly (weakly) prefers for a sincere student $i_{b}$ to become sophisticated if $i_{a}$ strictly (weakly) prefers her match in every Nash equilibrium of the multi-district school choice problem when $i_{b}$ is sophisticated to her match in every Nash equilibrium of the multi-district school choice problem when $i_{b}$ is sincere. We show that there can be many sophisticated students who strictly prefer for distinct sincere students to become sophisticated.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;3)$ model with category-probability lower bound $p$ and with one district using BM and the other using an arbitrary mechanism, there exists a set of sophisticated students of expected size at least $\tau n$ wherein each sophisticated student strictly prefers for a distinct sincere student to become sophisticated, and weakly prefers for all other sincere students to become sophisticated.
@@
Without loss of generality, let $R$ be the district that uses BM and let $L$ be the other district. We start by lower bounding the expected number of ordered quartets of students $(i_{1},i_{2},i_{3},i_{4})$ that satisfy the following conditions (as in the example from Section 3.1):
@@
    Conditions on sophistication types, constraint types, and residence:
@@
      Both $i_{3}$ and $i_{4}$ are sincere and reside in $R$.
@@
      $i_{4}$ most prefers a school $r_{1}\in R$ and second most prefers a school $r_{4}\in R$.
@@
      $i_{3}$ most prefers a school $r_{2}\in R$.
@@
      $i_{2}$ most prefers a school $\ell_{1}\in L$ and second most prefers a school $r_{3}\in R$.
@@
      $i_{1}$ has preference ordering $r_{2}\succ r_{1}\succ\ell_{1}$.
@@
      No student finds a school in $\{\ell_{1},r_{1},r_{2},r_{3},r_{4}\}$ acceptable except as above. [^11] That is: $i_{4}$ does not find $\ell_{1}$, $r_{2}$, or $r_{3}$ acceptable; $i_{3}$ does not find $\ell_{1}$, $r_{1}$, $r_{3}$, or $r_{4}$ acceptable; and so forth; and, no student other than $i_{1}$, $i_{2}$, $i_{3}$, and $i_{4}$ finds $\ell_{1},r_{1},r_{2},r_{3}$, or $r_{4}$ acceptable.
@@
      $i_{3}$ has priority over $i_{1}$ at $r_{2}$.
@@
As in the example in Section 3.1, under these conditions, $i_{2}$ strictly benefits from $i_{1}$ becoming sophisticated.
@@
We lower bound the probability that this set of conditions occurs for a specific $(i_{1},i_{2},i_{3},i_{4})$. Each of the three sets of conditions is independent. The conditions on sophistication types, constraint types, and residence are satisfied with probability at least $p^{4}$. The conditions on student preferences and school locations are satisfied with probability at least
$\left(\frac{1}{2}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-2}{n}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-3}{n}\cdot\frac{1}{2}\cdot\frac{n-4}{n-1}\cdot\frac{1}{2}\right)\cdot\\ \cdot\left(\frac{1}{n(n-1)(n-2)}\right)\cdot\left(\frac{n-5}{n}\cdot\frac{n-6}{n-1}\cdot\frac{n-7}{n-2}\right)^{n}.$
@@
In the expression above, each of the five parenthetical expressions corresponds to one of the five conditions on student preferences and school locations, and represents the condition’s probability (for the last expression: a lower bound on the condition’s probability), conditioned on the previous conditions. For sufficiently large nn, [^12] Note that $\lim_{n\rightarrow\infty}(1-\frac{r}{n-a})^{(n-b)}$, where $a$, $b$, and $r$ are constants, is $e^{-r}$. this probability can be lower bounded by [^13] The final $\nicefrac{{1}}{{2}}$ is to accommodate for all the multiplicands of the form $\frac{n-a}{n-b}$ or $(1-\frac{r}{n-a})^{(n-b)}$, which have a nonzero limit, not fully converging for finite large $n$.
$\frac{1}{2^{5}}\cdot\frac{(n-3)!}{n!}\cdot\frac{1}{e^{15}}\cdot\frac{1}{2}.$
@@
Finally, the conditions on school priorities are satisfied with probability $\nicefrac{{1}}{{2^{3}}}$.
@@
The number of ordered quartets of distinct students is $\tfrac{n!}{(n-4)!}$. List all such quartets and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{I_{j}}$ be the indicator random variable that represents whether the students in $I_{j}$ satisfy all of the above conditions. Then the expected number of quartets of students that satisfy all conditions is
$\mathbb{E}\left[\sum_{j=1}^{\tfrac{n!}{(n-4)!}}\mathbbm{1}_{I_{j}}\right]\geq\frac{n!}{(n-4)!}\cdot p^{4}\cdot\frac{1}{2^{3}}\cdot\frac{1}{2^{6}\cdot e^{15}}\cdot\frac{(n-3)!}{n!}=\frac{p^{4}(n-3)}{2^{9}\cdot e^{15}}.$
We observe that by construction, it is impossible for there to be two different quartets $(i_{1},i_{2},i_{3},i_{4})$ and $(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3},i^{\prime}_{4})$ that satisfy the above conditions such that $i^{\prime}_{1}=i_{1}$ or $i^{\prime}_{2}=i_{2}$. Assume for contradiction that for some $a\in\{1,2\}$, it is the case that $i^{\prime}_{a}=i_{a}$ for two such quartets. If $a=2$ then we note that $i^{\prime}_{1}=i_{1}$ as well, since $i_{1}$ is the only non-$i_{2}$ student who finds acceptable the school most preferred by $i_{2}$, and $i^{\prime}_{1}$ is the only non-$i^{\prime}_{2}$ student that finds acceptable the school most preferred by $i^{\prime}_{2}=i_{2}$. Therefore, regardless of the value of $a$, we have that $i^{\prime}_{1}=i_{1}$. However, this implies that $(i_{1},i_{2},i_{3},i_{4})=(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3},i^{\prime}_{4})$ because the identity of $i_{1}$ similarly uniquely determines those of $i_{2}$, $i_{3}$, and $i_{4}$, while the identity of $i^{\prime}_{1}=i_{1}$ uniquely determines those of $i^{\prime}_{2}$, $i^{\prime}_{3}$, and $i^{\prime}_{4}$ in the same way. This is a contradiction.
@@
To conclude, all such quartets of students involve distinct $i_{1}$” and $i_{2}$” students, so there are at least an expected $\frac{p^{4}}{2^{9}\cdot e^{15}}\cdot(n-3)$ sophisticated students who each strictly prefer for a distinct sincere student to become sophisticated. Also by construction, each such sophisticated student is unaffected by sincere students outside of her quartet, and the sophistication types of students $i_{3}$ and $i_{4}$ in the quartet do not affect the dynamic between students $i_{2}$ and $i_{1}$. [^14] This is because $i_{3}$ will always be assigned to her first choice school, and $i_{4}$ will either be assigned to $r_{1}$ if $i_{1}$ is sincere or $r_{4}$ if $i_{1}$ is sophisticated. Thus, these sophisticated students also weakly prefer for all other sincere students to become sophisticated. We choose $\tau=\frac{p^{4}}{2^{10}\cdot e^{15}}$ to satisfy the theorem statement. [^15] We made no attempt to optimize this value, as our goal is only to ascertain linearity in $n$ of the number of such students.
3.3 Necessity of a BM District
We conclude this section by proving that the condition in Theorem 3.1 of at least one district using BM is not an artifact of our proof, but rather dropping this condition invalidates the result.
Proposition 3.2.
For every multi-district school choice problem in which all districts use DA, there does not exist a sophisticated student and sincere student pair such that the sophisticated student strictly prefers for the sincere student to become sophisticated.
Proof.
Let there be a multi-district school choice problem with any number of districts, all of which use DA. Consider a sincere student $i$. We first show that $i$ enrolls in the same district whether she is sincere or sophisticated. This is straightforward if $i$ is constrained, as then she always enrolls in her district of residence $d(i)$.
Otherwise, suppose that $i$ is unconstrained. Let $d$ be the district that $i$ enrolls in when sincere. By the definition of sincere-unconstrained, $i$ weakly prefers enrolling in $d$ and truthfully ranking schools in $d$ to doing the same in any other district.
Note that because every district uses DA, if $i$ becomes sophisticated and enrolls in some district $d^{\prime}$, she still truthfully ranks the schools in $d^{\prime}$, thus obtaining the same outcome as if she enrolled in $d^{\prime}$ when sincere. Therefore, if $i$ becomes sophisticated, she again weakly prefers enrolling in $d$. If $i$ is assigned to any school when she enrolls in $d$ as a sincere student, then this preference is strict because $i$ has a strict preference ordering over schools. If $i$ is unassigned when she enrolls in $d$ as a sincere student, then she must be unassigned regardless of what district she enrolls in and what her sophistication level is, and $d=d(i)$. In either case, $i$ still enrolls in $d$ if she becomes sophisticated.
To conclude, $i$ behaves the same whether she is sincere or sophisticated: We showed that $i$ enrolls in the same district regardless of her sophistication level, and she reports her true preferences over schools in that district. As such, every student has the same outcome if $i$ becomes sophisticated. Since $i$ was arbitrary, it is therefore not possible for a sophisticated student to prefer for any sincere student to become sophisticated. ∎